OpenStudy (anonymous):

For 1 + 2 + 3 + ... + n = n(n + 1)/2, P(k + 1) would be A. 1 + 2 + 3 + ... + k = k(k + 1)/2 B. 1 + 2 + 3 + ... + (k + 1) = k(k+ 1)/2 C. 1 + 2 + 3 + ... + k = (k + 1)(k + 2)/2 D. 1 + 2 + 3 + ... + (k + 1) = (k + 1)(k + 2)/2

OpenStudy (rational):

1 + 2 + 3 + ... + n = n(n + 1)/2 simply replace `n` by `k+1`

OpenStudy (rational):

1 + 2 + 3 + ... + n = n(n + 1)/2 becomes : 1 + 2 + 3 + ... + (k+1) = (k+1)(k+1 + 1)/2

OpenStudy (anonymous):

THANK YOU SOOOO MUCH

OpenStudy (rational):

you're welcome !